OpenStudy (anonymous):

WILL MEDAL. given the first 4 terms and the number of terms of the geometeric series, find the finite sum. 2-10+50-250, n=8

OpenStudy (campbell_st):

well you have a 1st term of a = 2 what's the common ratio..?

OpenStudy (anonymous):

i got the answer of -132,208

OpenStudy (campbell_st):

now I need a calculator

OpenStudy (campbell_st):

ummm I think you have a calculation error

OpenStudy (solomonzelman):

2,-10,+50,-250, with 8 terms: 8 basically: \(\Large\color{blueviolet}{ \displaystyle \sum_{ k=1 }^{ 8 } 2(-5)^k}\)

OpenStudy (solomonzelman):

that is how I see the question

OpenStudy (campbell_st):

if you use the sum of a geometric series formula \[S_{n} = \frac{a(1 - r^n)}{1 - r}\] you know a = 2 and n = 8.... you need to find r. then substitute all 3 values to find the sum

OpenStudy (anonymous):

-130,208?

OpenStudy (campbell_st):

that looks a lot better